A large beaker is filled with ice cold water and is placed in a larger vessel containing a mixture of ice and water at zero degree temperature. A lump of aluminum weighing 120g initially cooled to -70C is then suspended inside the beaker. Explain why ice is formed on the aluminum lumps and not on the walls of the beaker.
Calculate the maximum weight of ice which can form on the aluminum = 9.10 x 10^2 Jkg^-1k^-1
1
Expert's answer
2020-02-17T09:24:17-0500

 The ice is formed on the aluminum lumps and not on the walls of the beaker because the initial temperature of aluminum is much lower than the ice.


"m_i L=m_a c_a (T-T_0)"

"m_i (3.33\cdot 10^5)=(0.12)(910) (0-(-70))"

"m_i=0.023\ kg=23\ g"